I need to select two odd numbers, two even numbers and zero for a formula that yields prime numbers...one such formula is x^2 -x + 41. Select some numbers for x, substitute them in a formula and see if prime numbers occur. Try to find a number for x that when substituted in the formula yields a composite number
Answer by ankor@dixie-net.com(22740)

I need to select two odd numbers, two even numbers and zero for a formula that
yields prime numbers...one such formula is x^2 -x + 41.
Select some numbers for x, substitute them in a formula and see if prime numbers occur.
:
Using the formula, find the prime numbers
x = 9: 9^2 - 9 + 41 = 113
x = 17: 17^2 - 17 + 41 = 313
x = 12: 12^2 - 12 + 41 = 173
x = 22: 22^2 - 22 + 41 = 503
x = 0: 0^2 - 0 + 41 = 41
All are prime numbers
:
Try to find a number for x that when substituted in the formula yields a composite number>
We know at least one value for x that will give a composite number, x=41